When ultraviolet light with a wavelength of 400nmfalls on a certain metal surface, the maximum kinetic energy of the emitted photoelectrons is 1.10eV .

What is the maximum kinetic energy K0 of the photoelectrons when light of wavelength 270nm falls on the same surface?
Answer #1

The law of external photoelectric effect (Einstein's equation) gives the sytem of equations:

\left\{\begin{matrix} \frac{hc}{\lambda _{1}}=L+E_{1}\\ \frac{hc}{\lambda _{2}}=L+E_{2}\\ \end{matrix}\right.

h = Planck's constant

c= speed of light

E1, E2 = kinetic energies

L = extraction work

Extract L from the 2 equations, equalize the terms (L is the same because is the same material that supports this process) and get E2.

[Take care at the units when make the numerical calculation.]